OpenStudy (anonymous):

@ganeshie8 Which equation could be used to solve the problem? Troy typed 375 word in 15 minutes. How many words per minute did he type (w)? A. w = 375 words • 15 minutes B. w = 375 words + 15 minutes C. w = 15 minutes ÷ 375 words D. w = 375 words ÷ 15 minutes

OpenStudy (anonymous):

D. w= 375/15

OpenStudy (anonymous):

show how u did that plz

OpenStudy (anonymous):

Just divide 375 by 15, tada!

OpenStudy (anonymous):

=25

OpenStudy (anonymous):

ok. so you have a total of 375 words in 15 min. you want to find how many are typed in one min. so you have 375 words/15 min = x words/1 min and you can do cross multiplication, you get 375=15x divide 375 by 15 and you get x or in this case w.

OpenStudy (anonymous):

hope that makes sense XD

OpenStudy (anonymous):

C

OpenStudy (anonymous):

if it was c then youd get 0.04 where as if you did 375 over 15 you'd get 25. which is more likely.

OpenStudy (anonymous):

is D.

OpenStudy (anonymous):

Yes. D is correct

OpenStudy (anonymous):

Which equation could be used to solve the problem? Oscar averaged 4 yards per carry in the football game. If he carried the ball 16 times, how many yards did he gain (y)? A. y = 4 • 16 B. y = 4 + 16 C. y = 4 ÷ 16 D. y = 16 ÷ 4

OpenStudy (anonymous):

In this case it'd be A. because the amount of times he carried the ball is greater than the original ratio you are given so if he carries the ball once, he averages 4 yards. if he carries the ball twice thats an average of 8 yards and so on and so forth, so A would get you the closest to what he would get after carrying the ball 16 times

OpenStudy (anonymous):

Ms. Huffman bought 15 cases of juice boxes for the town picnic. There were 24 juice boxes in each case. How many juice boxes did she buy?

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

ok so now youre going to get a fraction. divide 24 by 15 and that is your answer.

OpenStudy (anonymous):

oh sorry. i read that wrong. my bad

OpenStudy (anonymous):

24 in each case. 15 cases so multiply 24 by 15. you get 360 juice boxes